A total of $10,000 is deposited in two accounts. On one account the annual simple interest rate is 9.5%, on the other the annual simple interest is 13%. How much is invested in each account if the total annual interest earned is $1090.00
Answer by rfer(12677)

.095x+.13(10000-x)=1090
.095x+1300-.13x=1090
-.04x=-210
x=$5250 at 9.5%
10000-x=$4750 at 13%